OpenStudy (anonymous):

I've got an integration-by-parts monstrosity. I have the first part right (u-substitution), but when I have to use different variables, I mess up. Here's my problem: f(x) = k |x|^3 * e^(-x^2), where the integral from -infinity to infinity of the function should equal 1.

OpenStudy (anonymous):

I've gotten the u-sub as u = x^2 and du = 2x dx; so I can reduce this function down to 1= 2k [integral from 0 to b] x^3 * e^u du

OpenStudy (anonymous):

my attempt for integration by parts include w = u, dw = 1, dv = e^u, and v = e^u

OpenStudy (anonymous):

but I am just not having success. The answer, however, is k =1. I'm totally not getting that.

OpenStudy (anonymous):

anyone else getting stuck?

OpenStudy (anonymous):

My work:\[u = x^2; du = 2x;\] \[1= 2\int\limits_{0}^{b} kx^3 e^-x^2 dx\] \[1= 2k\int\limits\limits_{0}^{b} x^3 (e^(-x^2)) dx\]

OpenStudy (turingtest):

how about bringing the 2 back into the integral and using the sub \[dv=2xe^{-x^2}dx\]\[u=x^2\]I see interesting things happening I think, but I am multi-tasking

OpenStudy (anonymous):

\[1= k\int\limits\limits_{0}^{b} u* e^(u) du\]

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

now, w = u, dw = 1, dv = e^u, and v = e^u, right? So...

OpenStudy (anonymous):

Hello NoelGreco!

OpenStudy (anonymous):

\[\int\limits_{0}^{b}u*e^u = u*e^u -\int\limits_{0}^{b}(1)e^u du\]

OpenStudy (anonymous):

but how do I link that back to my original reduced eqn and solve for k?

OpenStudy (anonymous):

1= k[\[u*e^u - \int\limits_{0}^{b} e^u du\]] ???

OpenStudy (turingtest):

that's similar to what I'm getting

OpenStudy (anonymous):

but u* e^u - inifinity does NOT get k = 1. It gets madness.

OpenStudy (turingtest):

\[\int_{-\infty}^\infty k|x|^3e^{-x^2}dx=k\int_0^\infty x^2(2xe^{-x^2})dx\]\[u=x^2\implies du=2xdx\]\[dv=2xe^{-x^2}dx\implies v=-e^{-x^2}\]\[k\int_0^\infty x^2(2xe^{-x^2})dx=k\left(\cancel{-x^2e^{-x^2}|_0^\infty}^{\LARGE0}+\int2xe^{-x^2}dx\right)\]\[=k\left(-e^{-x^2}|_0^\infty\right)=k=1\]so it looks like k=1 :)

OpenStudy (anonymous):

woah. Hold on, I gotta see what I did wrong.

OpenStudy (turingtest):

I think the trick was in letting the 2 from the even integral trick into the integrand to make dv=2xe^(-x^2)

OpenStudy (anonymous):

this...looks so much easier than mine. Did you do integration by parts?

OpenStudy (anonymous):

OH

OpenStudy (turingtest):

sure did

OpenStudy (turingtest):

you actually gave me the idea, so it was a team effort :)

OpenStudy (anonymous):

I've gotta run for a minute but I'll be back online to make sure I get this. Thanks so much for your help (again, lol)! :D

OpenStudy (turingtest):

anytime, as usual :D